Consider the following condensed financial statements of Safe Money, Inc. The​ company's target rate of return is 20​%:

Safe Money, Inc.

Comparative Balance Sheet

As of December 31, 2018 and 2017

2018

2017

Assets

Cash

$82,000

$71,000

Accounts Receivable

61,600

27,500

Supplies

400

500

Property, Plant, and Equipment, net

299,000

197,000

Patents, net

158,000

103,000

Total Assets

$601,000

$399,000

Liabilities and Stockholders' Equity

Accounts Payable

$31,000

$33,000

Short-term Notes Payable

149,000

49,000

Long-term Notes Payable

199,000

129,000

Common Stock, no Par

201,000

168,500

Retained Earnings

21,000

19,500

Total Liabilities and Stockholders' Equity

$601,000

$399,000

Safe Money, Inc.

Income Statement

For the Year Ended December 31, 2018

Net Sales Revenue

$10,000,000

Cost of Goods Sold

5,900,000

Gross Profit

4,100,000

Operating Expenses

3,800,000

Operating Income

300,000

Other Income and (Expenses):

Interest Expense

(40,000)

Income Before Income Tax Expense

260,000

Income Tax Expense

91,000

Net Income

$169,000

Requirement 1. Calculate the​ company's ROI. Round all of your answers to four decimal places.

Operating income

/

Average total assets

=

ROI

/

=

  

%

Requirement 2. Calculate the​ company's profit margin ratio. Interpret your results.

Begin by selecting the formula to calculate profit margin​ ratio, and then enter the amounts to calculate the​ company's profit margin ratio. ​(Enter your final to the nearest hundredth of a​ percent, X.XX%)

Operating income

/

Net sales revenue

=

Profit margin ratio

/

=

%

The company's profit margin ratio tells us that they earn $

  

of

  

for every dollar of

.

Requirement 3. Calculate the​ company's asset turnover ratio. Interpret your results.

Begin by selecting the formula to calculate asset turnover​ ratio, and then enter the amounts to calculate the​ company's asset turnover ratio.

Net sales revenue

/

Average total assets

=

Asset turnover ratio

/

=

The company's asset turnover ratio explains how efficiently a company uses its

to generate

. The company is generating $

of

  

with every $1.00 of

.

  

Requirement 4. Use the expanded ROI formula to confirm your results from Requirement 1. Interpret your results.

​First, select the expanded ROI formula.

Profit margin ratio

x

Asset turnover ratio

=

ROI

Using the expanded formula, the ROI for Safe Money, Inc. is

  

%.

To determine what is driving a​ company's ROI, management often restates the ROI equation in its expanded form. The ROI calculated with the expanded formula

agrees

agrees

does not agree

with the ROI calculated in Requirement 1.

Requirement 5. Calculate the​ company's RI. Interpret your results.

​First, select the formula to calculate residual income​ (RI).

Operating income

(

Target rate of return

x

Average total assets

)

=

RI

The RI for Safe Money, Inc. is $

  

.

Safe Money comma Inc.Safe Money, Inc.

is

exceeding

not meeting

​management's target rate of return.

Choose from any list or enter any number in the input fields and then continue to the next question.
Answer #1

Safe Money inc.

R1. Return on investment = Operating income / Average total assets

= 300,000/(601,000+399,000)/2 = 300,000/(1000,000/2)

= 300,000/500,000 = 60%

Return on investment is 60 %

R2. Profit margin = Operating income / Net sales revenue

= 300,000/1000,000 = 30%

Profit margin ratio is 30%

R3. Assets turnover ratio = Net sales revenue / Average total assets

= 1000,000/500,000 = 0.5 times

Assets turnover ratio is 0.5 times

R4. Expanded ROI = Profit margin ratio × Assets turnover ratio

= 30% × 0.5 = 15%

Expanded ROI is 15%

R5. Residual income = Operating income - (Target rate of return × Average total assets)

= 300,000 - (20% × 500,000) = 300,000 - 150,000 = 150,000

Residual income is $ 150,000